OpenStudy (anonymous):

I need somebody to go step by step through factoring problems that have the leading coeeficiant infront of the first x 2x^2 + 11x + 12

OpenStudy (anonymous):

There are multiple ways to do this. How have you been taught?

OpenStudy (anonymous):

LA, i remember my teacher telling me i would have to multiply 2 and 12 so i get 24x but i do not remember further

OpenStudy (anonymous):

Okay. Once you get 24x, you are looking for factors that give you 24 and add up to 11.

OpenStudy (anonymous):

Your choices are: 1 * 24 2 * 12 3 * 8 4 * 6

OpenStudy (anonymous):

2x^2+8x+3x+12 2x(x+4)+3(x+4) (2x+3)(x+4)

OpenStudy (anonymous):

okay so 8 and 3

OpenStudy (anonymous):

and then u split the middle term or something right?

OpenStudy (anonymous):

Settling on 3 * 8, you rewrite 2x^2 + 11x + 12 as 2x^2 + 8x + 3x + 12 Notice we did not change the value of the equation. We did however go from 3 terms to 4 terms...

OpenStudy (anonymous):

ok let me write it down real wuick hang on

OpenStudy (anonymous):

Now you can factor by grouping.

OpenStudy (amistre64):

just write your factors with your 8 and 3; just remember to divide out the 2 you used in the beginning

OpenStudy (anonymous):

okay now do i group them into like two sets?

OpenStudy (anonymous):

one way is to make 3 columns and use trial and error 2x^2 +12 +11x ----------------------------- 2x * x 4 * 3 3 * 2x + 4 * x = 10x - incorrect 4 * 2 x + 3 * x = 11x - correct so factors are (2x + 3 )( x + 4 )

OpenStudy (anonymous):

2x^2 + 8x + 3x + 12 2x (x + 4) + 3 (x+4) Now again factor out what is common: (x + 4) (2x + 3)

OpenStudy (anonymous):

wait this is where i get confused

OpenStudy (anonymous):

i dont see wheere you are getting that

jimthompson5910 (jim_thompson5910):

See link below for full step by step solution http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.473453.html

OpenStudy (anonymous):

Okay. I split the four terms into two binomials: (2x^2 + 8x) + (3x + 12)

OpenStudy (amistre64):

then you might wanna try this way; you found 8 and 3 already so set it up: (x+8) (x+3) ; but we gotta divide out that 2 from the front (x+8/2) (x+3/2) ; thats better, now simplify (x+4) (x+3/2) ; if we still got a fraction, stick the bottom in front (x+4) (2x+3) and we are done

OpenStudy (anonymous):

I then factor each binomial. (2x^2 + 8x) + (3x + 12) Looking first for a GCF, I find 2x in the first binomial and 3 in the second binomial.

OpenStudy (anonymous):

okay i see it so far

OpenStudy (anonymous):

So (2x^2 + 8x) becomes 2x(x + 4) and (3x + 12) becomes 3(x+4)

OpenStudy (anonymous):

ok right there. where did 12 come from

OpenStudy (anonymous):

So that leaves me with: 2x(x + 4) + 3(x+4)

OpenStudy (anonymous):

alright i see it but how does 8x turn into just 4

OpenStudy (anonymous):

Recall we broke up (2x^2 + 8x) + (3x + 12) into 2 separate binomials.

OpenStudy (anonymous):

Because when we distribute 2x(x + 4) we go back to 2x^2 + 8x

OpenStudy (anonymous):

2x * x is 2x^2, 2x * 4 is 8x

OpenStudy (anonymous):

alright so this is the full problem, what one would be the answer? 2x2 + 11x + 12 Which of the following is one of the factors? (2x + 4) (x + 4) (x + 3) prime

OpenStudy (anonymous):

(x + 4)

jimthompson5910 (jim_thompson5910):

Since 2x^2+11x+12 = (x+4)(2x+3) as shown above, this means that x+4 and 2x+3 are both factors. So the answer is choice B since that's the only factor shown in the list of choices.